# A New Metric Function for SC-based Polar Decoders: Polarization, Pruning, and Fast Decoders

### v2

@@ Abstract
In this paper, we propose a method to obtain the optimal metric function at each depth of the polarization tree through a process we call polarization of the metric function. This polarization process generates an optimal metric at intermediate levels of the polarization tree, which can be applied in fast successive-cancellation-based (FSC) and SC list-based (FSCL) decoders—decoders that partially explore the binary tree representation. We prove that at each step of the polarization tree, the expected value of the metric function random variable is the mutual information of the corresponding channel, while its variance equals the varentropy of the channel—two parameters that are particularly relevant in finite block-length regimes. Additionally, we show that after polarization, the variances of the bit metrics approach zero for binary-input discrete memoryless channels (BI-DMCs). Moreover, we provide an estimate for calculating the variance of the binary-input additive white Gaussian noise (BI-AWGN) channel. We introduce a list-pruning strategy for FSCL decoding that retains only the paths whose metric values are close to the average. As a result, our method significantly reduces the number of required sorting operations in FSCL-based decoding algorithms. We also derive an upper bound, as a function of the polarized channel varentropy, on the probability that the distance between a bit-metric random variable and the bit-channel mutual information exceeds a given threshold. Leveraging this result, we further propose a varentropy-based list-pruning strategy for the SCL (VPSCL) decoding algorithm that adapts to the varentropy of the corresponding bit-channel.
@@ I Introduction
Polar codes, introduced by Arıkan [1], represent the first class of codes proven to achieve the capacity of memoryless symmetric channels with explicit constructions and low-complexity encoding and decoding. As an improvement to polar codes, polarization-adjusted convolutional (PAC) codes [2] are a new class of linear codes that demonstrate error-correction performance close to theoretical bounds in the short block-length regime [2, 3]. The main new ingredient of PAC codes is the use of convolutional codes as a pre-transformation that improves upon the weight spectrum of polar codes [4]. PAC codes are also equivalent to post-transforming polar codes with certain cyclic codes [5].
@@
Furthermore, leveraging our theoretical upper bounds, we propose a list-pruning strategy by specifying a threshold on the decoding metric to eliminate candidates from the list. The proposed threshold is a function of the polarized varentropy. A well-polarized channel has a smaller variance, thus, the absolute value of the list-pruning threshold can be smaller, enabling the SCL decoding algorithm to distinguish between correct and incorrect paths often without the need for sorting operations. This addresses another major bottleneck in the implementation of FSCL decoding, that is, the need to sort at each level of the polarization tree when the number of bifurcated paths exceeds $L$, in order to identify the best $L$ paths. For instance, for a high-rate PAC$(128,99)$ code constructed using the Reed-Muller rate profile, SCL decoding with a list of size $32$ delivers error correction performance on par with that of the Fano algorithm. In comparison, FSCL decoding only needs to visit $28$ nodes of the polarization tree, a substantial reduction from the $254$ nodes needed for conventional SCL decoding. Furthermore, our approach cuts the number of sorting operations of FSCL algorithm to about $33\%$ of the previous requirement, which further enhances latency performance at a high $E_{b}/N_{0}$ regime. For a PAC$(1024,512)$ code constructed using the polar rate profile, the reduction in sorting operations for the FSCL algorithm is even more substantial, decreasing to about $5\%$ of the previous requirement.
@@
As another application, the new decoding metric can be used to bridge the theoretical analyses of convolutional codes (CCs) to polar and PAC codes, providing a new perspective for studying these codes. In convolutional codes (CCs), the decoder employs a tree search algorithm and utilizes an optimal metric function to determine the path that is closest to the transmitted codeword. Many theoretical analyses of CCs rely on the assumption that the bit metric function variables are independent and identically distributed (i.i.d.) for each channel use [16, Ch. 6], [17, Ch. 6]. In a recent work, inspired by the theoretical analyses of CC codes, the authors have leveraged the properties of the new decoding metric to upper bound the computation distribution of polar-like codes [18].
@@ II-B Polar and Reed-Muller Codes
As a well-studied family of linear block codes, Reed-Muller (RM) codes benefit from their specific and universal construction (i.e., not channel dependent), favorable weight spectrum, and appealing structural properties [19, 20]. RM codes can be also viewed as codes in the general family of polar codes. More specifically, an RM$(r,m)$ code, with length $N=2^{m}$ and dimension $K=\binom{m}{0}+\binom{m}{1}+\cdots+\binom{m}{r}$, where $\binom{m}{t}$ is a binomial coefficient, can be constructed by selecting all $K$ rows of the matrix $\mathbf{F}_{N}$ that have Hamming weights greater than or equal to $2^{m-r}$. The selection of row indices is a key distinction between RM codes and that of polar codes driven by their successive cancellation (SC)-type decoders. Additionally, the code dimension $K$ in an RM code can take on $m+1$ distinct values, whereas in polar codes $K$ can be arbitrarily set as long as $1\leq K\leq N$. In this paper, we use the notation RM$(N,K)$ instead of the more conventional RM$(r,m)$ notation.
@@ III Metric Polarization
[…]
@@
[…]
@@
Continuing with the recursive process, Fig. 3 illustrates the two-step polarization. The metric function ϕsuperscriptitalic-$\phi^{--}$ corresponding to channel $W^{--}$ is defined as
@@
Similarly, for channels $W^{-+}$, $W^{+-}$, and $W^{++}$, we define the metric functions ϕ+superscriptitalic-$\phi^{-+}$, ϕ+superscriptitalic-$\phi^{+-}$, and ϕ++superscriptitalic-$\phi^{++}$. Assuming the inputs of the synthesized channels are uniformly distributed, an optimal decoder should maximize the value of these metric functions.
@@
ϕsuperscriptitalic-$\phi^{--}$ϕ+superscriptitalic-$\phi^{-+}$ϕ+superscriptitalic-$\phi^{+-}$ϕ++superscriptitalic-$\phi^{++}$--$+$ (a) 111100 (b)
Figure 4: (a) Polarization and (b) code tree of a PAC$(4,2)$ code.
Fig. 4 illustrates the polarization tree and code tree corresponding to a PAC$(4,2)$ code, where $\mathcal{A}=\{2,3\}$. For $K$ data bits, the code tree contains $2^{K}$ distinct paths from the root to each of the leaf nodes. The decoder’s task is to identify the path that corresponds to the transmitted data. This involves determining whether to proceed up (corresponding to data bit 00) or down (corresponding to data bit $1$). The polarization tree has at most $\log_{2}(N)$ levels. Using the polarization tree, the decoder computes the bit metric function values and subsequently traverses the code tree. Each leaf node of the polarization tree corresponds to a specific level of the code tree. For a Fast SCL decoder, it is not necessary to traverse the polarization tree up to the leaf level to determine the correct branch in the code tree. The following proposition characterizes the polarization of the metric function. As discussed in [16, p. 206], we employ the standard ensemble of random linear codes for convolutional codes to conduct our analysis. This ensemble can be represented by the pair $(\mathbf{T},\mathbf{c})$ in the form $\mathbf{v}\mathbf{T}+\mathbf{c}$, where $\mathbf{c}$ is an arbitrary, fixed binary vector of length $N$.
@@
$\begin{split}&\mathbb{E}_{\prescript{}{i}{R}_{j},(\prescript{}{i}{\mathbf{Y}},\prescript{}{i}{\mathbf{R}}_{1}^{j-1})}\left[\phi^{\{-,+\}^{k}}(\prescript{}{i}{R}_{j};\prescript{}{i}{\mathbf{Y}},\prescript{}{i}{\mathbf{R}}_{1}^{j-1})\right]\\ &=\sum_{\prescript{}{i}{r}_{j}}P(\prescript{}{i}{r}_{j})\sum_{(\prescript{}{i}{\mathbf{y}},\prescript{}{i}{\mathbf{r}}_{1}^{j-1})}P(\prescript{}{i}{\mathbf{y}},\prescript{}{i}{\mathbf{r}}_{1}^{j-1}\mid\prescript{}{i}{r}_{j})\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\phi^{\{-,+\}^{k}}(\prescript{}{i}{r}_{j};\prescript{}{i}{\mathbf{y}},\prescript{}{i}{\mathbf{r}}_{1}^{j-1})\\ &=\sum_{\prescript{}{i}{r}_{j}}\sum_{(\prescript{}{i}{\mathbf{y}},\prescript{}{i}{\mathbf{r}}_{1}^{j-1})}P(\prescript{}{i}{r}_{j})P(\prescript{}{i}{\mathbf{y}},\prescript{}{i}{\mathbf{r}}_{1}^{j-1}\mid\prescript{}{i}{r}_{j})\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\log_{2}\left(\frac{P(\prescript{}{i}{\mathbf{y}},\prescript{}{i}{\mathbf{r}}_{1}^{j-1}\mid\prescript{}{i}{r}_{j})}{P(\prescript{}{i}{\mathbf{y}},\prescript{}{i}{\mathbf{r}}_{1}^{j-1})}\right)\\ &=I\left(W^{\{-,+\}^{k}}(\prescript{}{i}{r}_{j};\prescript{}{i}{\mathbf{y}},\prescript{}{i}{\mathbf{r}}_{1}^{j-1})\right).\\ \end{split}$ (20)
@@
[…]
@@
[…]
@@
[…]
@@
Fig. 5 plots $K(t)$ from (27) and compares it with our proposed approximation in (28). Fig. 6 also plots $\text{Var}_{X,Y}\left[\phi(X,Y)\right]$ using $K(t)$ and $J(t)$ and compares it with the variance approximation obtained from $J_{\text{approx}}(t)$ and $K_{\text{approx}}(t)$ functions. Additionally, we plot the variance of the channel approximation obtained using the method introduced in [29] for computing channel capacity. As shown in this figure, to keep the variance of the metric values close to zero, $E_{b}/N_{0}$ should be greater than 101010~{}10dB. At 2.52.52.5~{}2.5dB, the variance exceeds $0.5$. In theorem 2, we prove that for polar and PAC codes, the variance of the bit-metric values almost surely approaches zero. For $N=1024$, Fig. 7 shows the sorted polarized bit-channel capacities and the corresponding variances of a BI-AWGN channel at $E_{b}/N_{0}=2.5$ dB.
For the case of a BEC($\epsilon$) with the erasure probability of $\epsilon$, the variance is $V=\epsilon(1-\epsilon)$. For $\epsilon=0.3$, the metric variance is $V=0.21$. Note that the maximum value of $V$ achieves at $\epsilon=0.5$, where it attains $V=0.25$. For $N=1024$, Fig. 8 shows the sorted polarized bit-channel capacities and the corresponding variances of BEC$(0.3)$. The following theorem proves the polarization effect for the variance of the polarized metric function values of BECs, and we later generalize this to any B-DMC.
@@
Figure 7: Plot of sorted $I(W_{N}^{(i)})$ and corresponding $V(W_{N}^{(i)})$ versus channel indices $10241,2,\cdots,N=$ 1024 for a BI-AWGN at $E_{b}/N_{0}=2.5$ dB.
Figure 8: Plot of sorted $I(W_{N}^{(i)})$ and corresponding $V(W_{N}^{(i)})$ versus channel indices $10241,2,\cdots,N=$ 1024 for a BEC with $\epsilon=0.3$.
@@
If the crossover probability $\delta$ approaches $0.5$ (i.e., when the channel is highly noisy), the capacity tends to zero. On the other hand, as $\delta$ approaches 00 or $1$ (i.e., for a noiseless channel), the capacity approaches 1. In each of these limiting cases, the channel dispersion $V$ tends to zero.
@@
0.79440.79440.79440.79440.63860.63860.63860.63860.41980.41980.41980.41980.1891Rate-00.65040.65040.65040.65040.4349Rate-00.8663SPC0.85760.85760.85760.85760.74060.74060.74060.74060.5579Rate-00.9238SPC0.9753SPC0.95080.95080.95080.95080.90540.90540.90540.90540.8230Type-I0.98850.98850.98850.98850.9774REP0.9998Rate-10.99660.99660.99660.99660.99330.99330.99330.99330.9868REP0.9999SPC1Rate-1
Figure 9: Tree diagram of the average value of the partially polarized metric functions corresponding to the polarization tree of Fast SCL decoding of PAC$(64,32)$ code using MC rate profile.

### v3

@@ Abstract
In this paper, we propose a method to obtain the optimal metric function at each depth of the polarization tree through a process we call polarization of the metric function. This polarization process generates an optimal metric at intermediate levels of the polarization tree, which can be applied in fast successive-cancellation-based (FSC) and SC list-based (FSCL) decoders—decoders that partially explore the binary tree representation. We prove that at each step of the polarization tree, the expected value of the metric function random variable is the mutual information of the corresponding channel, while its variance equals the varentropy of the channel—two parameters that are particularly relevant in finite block-length regimes. Additionally, we show that after polarization, the variances of the bit metrics approach zero for binary-input discrete memoryless channels (BI-DMCs). Moreover, we provide an estimate for calculating the variance of the binary-input additive white Gaussian noise (BI-AWGN) channel. We introduce a list-pruning strategy for FSCL decoding that retains only the paths whose metric values are close to the average. As a result, our method significantly reduces the number of required sorting operations in FSCL-based decoding algorithms. We also derive an upper bound, as a function of the polarized channel varentropy, on the probability that the distance between a bit-metric random variable and the bit-channel mutual information exceeds a given threshold. Leveraging this result, we further propose a varentropy-based list-pruning strategy for the SCL (VPSCL) decoding algorithm that adapts to the varentropy of the corresponding bit-channel. Our proposed pruning strategy also benefits stack decoding (VPStack) by discarding partial paths and avoiding unnecessary extensions.
@@ I Introduction
Polar codes, introduced by Arıkan [1], represent the first class of codes proven to achieve the capacity of memoryless symmetric channels with explicit constructions and low-complexity encoding and decoding. As an improvement to polar codes, polarization-adjusted convolutional (PAC) codes [2] are a new class of linear codes that demonstrate error-correction performance close to theoretical bounds in the short block-length regime [2, 3]. The main new ingredient of PAC codes is the use of convolutional codes as a pre-transformation that improves upon the weight spectrum of polar codes [4, 5]. PAC codes are also equivalent to post-transforming polar codes with certain cyclic codes [5].
@@
Furthermore, leveraging our theoretical upper bounds, we propose a list-pruning strategy by specifying a threshold on the decoding metric to eliminate candidates from the list. The proposed threshold is a function of the polarized varentropy. A well-polarized channel has a smaller variance, thus, the absolute value of the list-pruning threshold can be smaller, enabling the SCL decoding algorithm to distinguish between correct and incorrect paths often without the need for sorting operations. This addresses another major bottleneck in the implementation of FSCL decoding, that is, the need to sort at each level of the polarization tree when the number of bifurcated paths exceeds $L$, in order to identify the best $L$ paths. For instance, for a high-rate PAC(128,99)(128,99) code constructed using the Reed-Muller rate profile, SCL decoding with a list of size $32$ delivers error correction performance on par with that of the Fano algorithm. In comparison, FSCL decoding requires visiting only $28$ nodes of the polarization tree—each corresponding to a time step—which represents a substantial reduction from the $254$ time steps required by conventional SCL decoding. Furthermore, our approach cuts the number of sorting operations of FSCL algorithm to about $33\%$ of the previous requirement, which further enhances latency performance at a high $E_{b}/N_{0}$ regime. For a PAC(1024,512)(1024,512) code constructed using the polar rate profile, the reduction in sorting operations for the FSCL algorithm is even more substantial, decreasing to about $5\%$ of the previous requirement.
Stack decoding [16, 17] is a sequential decoding algorithm that explores each path on the code tree at most once and requires sorting the paths in the stack at each decoding step. We have extended our pruning strategy to stack decoding (referred to as VPStack) to enable early discarding of unlikely paths during decoding. Numerical results for PAC(128,64)(128,64) demonstrate that VPStack decoder achieves error-correction performance comparable to that of the conventional stack decoder. Moreover, at $E_{b}/N_{0}=3.5$ dB, VPStack reduces the number of paths stored in the stack to approximately 12% of those in conventional stack decoding.
@@
As another application, the new decoding metric can be used to bridge the theoretical analyses of convolutional codes (CCs) to polar and PAC codes, providing a new perspective for studying these codes. In convolutional codes (CCs), the decoder employs a tree search algorithm and utilizes an optimal metric function to determine the path that is closest to the transmitted codeword. Many theoretical analyses of CCs rely on the assumption that the bit metric function variables are independent and identically distributed (i.i.d.) for each channel use [18, Ch. 6], [19, Ch. 6]. In a recent work, inspired by the theoretical analyses of CC codes, the authors have leveraged the properties of the new decoding metric to upper bound the computation distribution of polar-like codes [20].
The main contributions of this work are as follows.
  • We introduced the polarization of the metric function and showed that the resulting metric random variables become independent and identically distributed (i.i.d.) across the polarized channels.
  • We established a novel connection between the decoding algorithm’s metric function and the channel’s varentropy.
  • We proposed an efficient approximation method for computing the varentropy of the binary-input AWGN (BI-AWGN) channel.
  • We developed a varentropy-based list pruning technique for both SCL (VPSCL) and stack (VPStack) decoding algorithms, which effectively identifies and discards incorrect paths to reduce computational complexity.
@@ II-B Polar and Reed-Muller Codes
As a well-studied family of linear block codes, Reed-Muller (RM) codes benefit from their specific and universal construction (i.e., not channel dependent), favorable weight spectrum, and appealing structural properties [21, 22]. RM codes can be also viewed as codes in the general family of polar codes. More specifically, an RM(r,m)(r,m) code, with length $N=2^{m}$ and dimension $K=\binom{m}{0}+\binom{m}{1}+\cdots+\binom{m}{r}$, where $\binom{m}{t}$ is a binomial coefficient, can be constructed by selecting all $K$ rows of the matrix $\mathbf{F}_{N}$ that have Hamming weights greater than or equal to $2^{m-r}$. The selection of row indices is a key distinction between RM codes and that of polar codes driven by their successive cancellation (SC)-type decoders. Additionally, the code dimension $K$ in an RM code can take on $m+1$ distinct values, whereas in polar codes $K$ can be arbitrarily set as long as $1\leq K\leq N$. In this paper, we use the notation RM(N,K)(N,K) instead of the more conventional RM(r,m)(r,m) notation.
@@ II-D FSC, FSCL and Metric Functions
II-E Stack decoding
Stack decoding is a sequential decoding algorithm that explores the code tree by maintaining all explored paths in a stack [16, 17]. The algorithm aims to identify the path on the code tree that corresponds to the transmitted codeword, starting at the root of the tree with an initial path metric of zero. The decoder continues exploring the tree until it reaches a leaf node or satisfies a predefined stopping criterion. At each iteration, the decoder selects (pops) the path with the highest metric from the stack. If the current bit is an information bit, the algorithm bifurcates the path, computes the metrics of both successor paths, and pushes them back into the stack. If the bit is frozen, only the single valid successor path is extended, and its metric is computed and stored. For a memory-efficient implementation of stack decoding for PAC codes, we refer the reader to [30, Ch. 6].
@@ III Metric Polarization
$\mathbf{s}_{1}^{N/2~~~}$$\mathbf{s}_{N/2+1}^{N}$$\mathbf{x}_{1}^{N/2~~~}$$\mathbf{x}_{N/2+1}^{N}$$W^{N/2}$$W^{N/2}$$\mathbf{y}_{1}^{N/2~~~}$$\mathbf{y}_{N/2+1}^{N}$
@@
$\mathbf{v}_{1}^{N/4~~~~}$$\mathbf{v}_{N/4+1~}^{N/2}$$\mathbf{v}_{N/2+1~}^{3N/4}$$\mathbf{v}_{3N/4+1}^{N}$$\mathbf{s}_{1}^{N/4~~~~}$$\mathbf{x}_{1}^{N/4~~~~}$$\mathbf{s}_{N/4+1~}^{N/2}$$\mathbf{x}_{N/4+1~}^{N/2}$$\mathbf{s}_{N/2+1~}^{3N/4}$$\mathbf{x}_{N/2+1~}^{3N/4}$$\mathbf{s}_{3N/4+1}^{N}$$\mathbf{x}_{3N/4+1}^{N}$$W^{N/4}$$W^{N/4}$$W^{N/4}$$W^{N/4}$$\mathbf{y}_{1}^{N/4~~~}$$\mathbf{y}_{N/4+1~}^{N/2}$$\mathbf{y}_{N/2+1~}^{3N/4}$$\mathbf{y}_{3N/4+1}^{N}$
@@
Continuing with the recursive process, Fig. 3 illustrates the two-step polarization. The metric function $\phi^{--}$ corresponding to channel $W^{--}$ is defined as
@@
Similarly, for channels $^{-+}, W+W^{+-}$, and $W^{++}$, we define the metric functions $\phi^{-+}$, $\phi^{+-}$, and $\phi^{++}$. Assuming the inputs of the synthesized channels are uniformly distributed, an optimal decoder should maximize the value of these metric functions.
@@
$\phi^{--}$$\phi^{-+}$$\phi^{+-}$$\phi^{++}$-++ (a) 110 (b)
Figure 4: (a) Polarization and (b) code tree of a PAC(4,2)(4,2) code.
Fig. 4 illustrates the polarization tree and code tree corresponding to a PAC(4,2)(4,2) code, where $\mathcal{A}=\{2,3\}$. For $K$ data bits, the code tree contains $2^{K}$ distinct paths from the root to each of the leaf nodes. The decoder’s task is to identify the path that corresponds to the transmitted data. This involves determining whether to proceed up (corresponding to data bit 0) or down (corresponding to data bit $1$). The polarization tree has at most $\log_{2}(N)$ levels. Using the polarization tree, the decoder computes the bit metric function values and subsequently traverses the code tree. Each leaf node of the polarization tree corresponds to a specific level of the code tree. For a Fast SCL decoder, it is not necessary to traverse the polarization tree up to the leaf level to determine the correct branch in the code tree. The following proposition characterizes the polarization of the metric function. As discussed in [18, p. 206], we employ the standard ensemble of random linear codes for convolutional codes to conduct our analysis. This ensemble can be represented by the pair $(\mathbf{T},\mathbf{c})$ in the form $\mathbf{v}\mathbf{T}+\mathbf{c}$, where $\mathbf{c}$ is an arbitrary, fixed binary vector of length $N$.
@@
[…]
@@
$J(t)\triangleq 1-\frac{1}{t\sqrt{2\pi}}\int_{-\infty}^{+\infty}e^{\frac{-(u-\frac{t^{2}}{2})^{2}}{2t^{2}}}\log_{2}(1+e^{-u})\,du,$ (23)
@@
$\begin{split}&\text{Var}_{X,Y}\left[\phi(X,Y)\right]\\ &=\sum_{x\in\{-1,1\}}\int_{-\infty}^{+\infty}\frac{1}{2}p(y|x)\phi^{2}(X,Y)dy-I(W)^{2}.\end{split}$ (25)
@@
$K(t)\triangleq 1-\frac{1}{t\sqrt{2\pi}}\int_{-\infty}^{+\infty}e^{-\frac{(u-\frac{t^{2}}{2})^{2}}{2t^{2}}}\log_{2}^{2}\left(1+e^{-u}\right)dy.$ (27)
@@
Fig. 5 plots $K(t)$ from (27) and compares it with our proposed approximation in (28). Fig. 6 also plots $\text{Var}_{X,Y}\left[\phi(X,Y)\right]$ using $K(t)$ and $J(t)$ and compares it with the variance approximation obtained from $J_{\text{approx}}(t)$ and $K_{\text{approx}}(t)$ functions. Additionally, we plot the variance of the channel approximation obtained using the method introduced in [32] for computing channel capacity. As shown in this figure, to keep the variance of the metric values close to zero, $E_{b}/N_{0}$ should be greater than 1010~dB. At 2.52.5~dB, the variance exceeds $0.5$. In theorem 2, we prove that for polar and PAC codes, the variance of the bit-metric values almost surely approaches zero. For $N=1024$, Fig. 7 shows the sorted polarized bit-channel capacities and the corresponding variances of a BI-AWGN channel at $E_{b}/N_{0}=2.5$ dB.
For the case of a BEC($\epsilon$) with the erasure probability of $\epsilon$, the variance is $V=\epsilon(1-\epsilon)$. For $\epsilon=0.3$, the metric variance is $V=0.21$. Note that the maximum value of $V$ achieves at $\epsilon=0.5$, where it attains $V=0.25$. For $N=1024$, Fig. 8 shows the sorted polarized bit-channel capacities and the corresponding variances of BEC(0.3)(0.3). The following theorem proves the polarization effect for the variance of the polarized metric function values of BECs, and we later generalize this to any B-DMC.
@@
Figure 7: Plot of sorted $I(W_{N}^{(i)})$ and corresponding $V(W_{N}^{(i)})$ versus channel indices $1,2,\cdots,N=1024$ for a BI-AWGN at $E_{b}/N_{0}=2.5$ dB.
Figure 8: Plot of sorted $I(W_{N}^{(i)})$ and corresponding $V(W_{N}^{(i)})$ versus channel indices $1,2,\cdots,N=1024$ for a BEC with $\epsilon=0.3$.
@@
If the crossover probability $\delta$ approaches $0.5$ (i.e., when the channel is highly noisy), the capacity tends to zero. On the other hand, as $\delta$ approaches 0 or $1$ (i.e., for a noiseless channel), the capacity approaches 1. In each of these limiting cases, the channel dispersion $V$ tends to zero.
@@
0.79440.79440.63860.63860.41980.41980.1891Rate-00.65040.65040.4349Rate-00.8663SPC0.85760.85760.74060.74060.5579Rate-00.9238SPC0.9753SPC0.95080.95080.90540.90540.8230Type-I0.98850.98850.9774REP0.9998Rate-10.99660.99660.99330.99330.9868REP0.9999SPC1Rate-1
Figure 9: Tree diagram of the average value of the partially polarized metric functions corresponding to the polarization tree of Fast SCL decoding of PAC(64,32)(64,32) code using MC rate profile.